OpenStudy (anonymous):

Write the equation in standard form 45y^2 - 320x^2 + 6 = 2886. Write the standard form of 9x^2 + 36x + 25y^2 - 50y = 164

OpenStudy (jack1):

Standard form, is that the one that's Ax + By = C ??? or is it y = mx + c ???

OpenStudy (anonymous):

Ax + By = C

OpenStudy (jack1):

coolski, so ur almost there already with the 1st won \(\large 45y^2 - 320x^2 + 6 = 2886\) so minus 6 from both sides of the equals sign, what r u left with?

OpenStudy (anonymous):

45y^2 - 320x^2 = 2880

OpenStudy (anonymous):

none of the conic section's standards look like that

OpenStudy (jack1):

cool, so now everything is divisible by 5 so divide everything by 5

OpenStudy (anonymous):

9y^2 + 64x^2 =576

OpenStudy (anonymous):

y isn't supposed to be by itself.

OpenStudy (jack1):

9y^2 - 64x^2 =576